OpenStudy (anonymous):

simplify 4 radical 98

OpenStudy (anonymous):

hint... 98 = 2 * 49

OpenStudy (anonymous):

yea ! And we have : 49=7^2 can say : 4 radical 98 = 7 * 4 * radical 2 = 28 radical 2 got it ??:)

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

ya but what does it mean when it has the little 4

OpenStudy (anonymous):

that's the 4th root. \[\sqrt[4]{81}=\sqrt[4]{3^4}=3\]

OpenStudy (anonymous):

hint 112 = 7 * 16

OpenStudy (anonymous):

Okay... Um...